2. A high-speed electron having a mass of 9 Γ— 10βˆ’31π‘˜π‘” is moving at right angle to 0.75-T magnetic field and has a speed of 2.5 Γ— 107π‘š/𝑠. What is the size of the force acts on the high-speed electron? What is the magnitude of acceleration of the particle?

1
Expert's answer
2021-10-31T18:15:53-0400

"B=0.75 \;T \\

v=2.5 \times 10^7 \;m/s \\

m = 9 \times 10^{-31} \;kg \\

q = e = 1.6 \times 10^{-19} \;C \\

θ=90°"

Magnetic force

"F = qvBsinθ \\

F = 1.6 \times 10^{-19} \times 2.5 \times 10^7 \times 0.75 \times sin 90° \\

F = 3 \times 10^{-12} \;N"

Acceleration

"a = \frac{F}{m} \\

a = \frac{3 \times 10^{-12}}{9 \times 10^{-31}} \\

a = 0.333 \times 10^{19} = 3.33 \times 10^{18} \;m/s^2"